OpenStudy (zzr0ck3r):

From knowing \[7c=\sum_{i-1}^{50-c}k_i\] and \(c\choose 2 \)=\( \sum_{i=1}^{50-c}\)\(k_i\choose 2 \) how can I get to \[\sum_{i=1}^{50-c}(k_i-\mu)^2=(50-c)\mu^2-14c\mu+c^2+6c\] for some arbitrary \(\mu\)?

OpenStudy (zzr0ck3r):

The book I am reading has no explination, it just says "from the last two equations, it follows that for any real number \(\mu\)..." and then gives that last equality.

hartnn (hartnn):

you just need \(\sum k_i^2 \) to be = c^2 +6c, right? the rest is trivial

OpenStudy (zzr0ck3r):

it is late, I dont see the trivial part lol. But I have been staring at this for hours....

hartnn (hartnn):

\(\sum \dfrac{k_i(k_i-1)}{2} = \dfrac{c(c-1)}{2}\)

hartnn (hartnn):

ohh lol \(\sum k_i^2-2\sum k_i\mu +\sum \mu^2 \)

OpenStudy (zzr0ck3r):

I didn't even remember we can do that, but sure it makes sense....

hartnn (hartnn):

\(\mu \) is a constant so, \(\sum k_i^2-2\mu \sum k_i +\sum \mu^2 =\sum k_i^2-2\mu(7c)+\mu^2(50-c) \)

hartnn (hartnn):

so coming back to \(\sum k_i^2\) \(\sum \dfrac{k_i(k_i-1)}{2} = \dfrac{c(c-1)}{2}\) \(\sum k_i^2 -\sum k_i = c^2 -c \\ \sum k_i^2 = c^2-c+7c = ...\) just some algebraic manipulations:)

OpenStudy (zzr0ck3r):

wow

OpenStudy (zzr0ck3r):

you are the man

OpenStudy (zzr0ck3r):

I feel silly, but I'll take it :)

OpenStudy (zzr0ck3r):

this is for a 600 lvl graph theory class, and I can't to the algebra from precalculus ...

hartnn (hartnn):

its late, you'd surely have got it in the morning :P

OpenStudy (rsadhvika):

is there a way to work this combinatorially ?

OpenStudy (zzr0ck3r):

yeah, let's go with that.

OpenStudy (zzr0ck3r):

or right when I go to sleep....

OpenStudy (zzr0ck3r):

I have done enough counting for the night @rsadhvika but dig in :)

OpenStudy (rsadhvika):

my browser is not diplaying latex

OpenStudy (rsadhvika):

\(\sum\limits_{i=1}^n \dbinom{k_i}{2}\) is the total number of ways of choosing \(2n\) people from \(n\) different groups, picking \(2\) people from each group.

OpenStudy (rsadhvika):

\(\sum\limits_{i=1}^{n}k_i\) is the total number of people in all the groups.

OpenStudy (rsadhvika):

im using \(50-c=n\) for simplicity..

OpenStudy (rsadhvika):

\(\mu\) maybe number of ppl that you don't want to pick who have bad record maybe..

OpenStudy (rsadhvika):

but \(\mu\) is fixed across all the groups

OpenStudy (rsadhvika):

that means you don't want to pick \(\mu\) number of people from each group

OpenStudy (rsadhvika):

or maybe you're simply calculating RMS value of the observations and this may not have any useful meaning combinatorially..